# Out-of-Distribution Generalization via Invariant Trajectories for Multimodal Large Language Model Editing

### v1

@@ Abstract
Knowledge editing emerges as a crucial technique for efficiently correcting incorrect or outdated knowledge in large language models (LLMs). Existing editing methods for unimodal LLM rely on a rigid parameter-to-output mapping, which causes causal-underfit and causal-overfit in cascaded reasoning for Multimodal LLM (MLLM). In this paper, we reformulate MLLM editing as an out-of-distribution (OOD) generalization problem, where the goal is to discern semantic shift with factual shift and thus achieve robust editing among diverse cross-modal prompting. The key challenge of this OOD problem lies in identifying invariant causal trajectories that generalize accurately while suppressing spurious correlations. To address it, we propose ODEdit,a plug-and-play invariant learning based framework that optimizes the tripartite OOD risk objective to simultaneously enhance editing reliability, locality, and generality. We further introduce an edit trajectory invariant learning method, which integrates a total variation penalty into the risk minimization objective to stabilize edit trajectories against environmental variations. Theoretical analysis and extensive experiments demonstrate the effectiveness of ODEdit.
@@ 1 Introduction
With rapid applications of large language models (LLM) (Liu et al., 2024, 2025; Su et al., 2025), ensuring their knowledge correctness and currency in a cost-efficient manner has become a critical concern. Knowledge editing (De Cao et al., 2021; Wang et al., 2023, 2024b) is an emerging technique that supports data-efficient modifications on pre-trained models within a specific scope of knowledge. Existing methods have two categories, i.e., , i) parameter-adjusting (Meng et al., 2022b; Fang et al., 2024; Jiang et al., 2025) directly tune a subset of parameters in the original model, and ii) model-extending (Huang et al., 2023; Hartvigsen et al., 2023; Yu et al., 2024) attaches auxiliary components while keeping the backbone parameters intact. A unifying goal is to promote the precision (Reliability) and generalization (Generality) of LLM perception on the editing knowledge, without compromising irrelevant knowledge outside the editing scope (Locality).
Figure 1: The motivation of ODEdit. The left presents why previous editing work targeted at unimodal LLM is not effective in MLLM. The right denotes two shifts in this editing OOD problem.
Despite these advances, current studies remain largely confined to unimodal LLMs, leaving open their extension to multimodal LLMs (MLLMs) (Guo et al., 2025; Luo et al., 2025a; Zhou et al., 2025b). As Figure 1(a) illustrates, both parameter-adjusting and model-extending methods operationalize editing as a rigid mapping from parameter modifications $\Delta W$ or auxiliary component modifications $\Delta M$ to output variations $\Delta Y$, distilled from a limited training cases. However, in the Structural Causal Model (SCM) view (Li et al., 2024c; Zhou et al., 2024), the forward computation graph of an MLLM is a structural causal model: each module implements a structural equation, forming a directed causal chain as unimodal perception → cross-modal alignment → shared semantic reasoning. Any local change to a module or parameter inevitably propagates downstream and alters subsequent states, and their effects are mediated by all later causal mechanisms. Consequently, rigid mapping in MLLM editing easily induces two issues:
• Causal Underfit. It fails to disentangle the coherent shared causal structures that span diverse cross-modal contexts. This precludes MLLM from discovering reusable semantic substrates, reducing editing to a brittle case-wise alignment exercise rather than principled semantic generality.
• Causal Overfit. Such a mapping drives MLLM to memorize entrenched linkages between local features and outputs. The inflexible causal chains misguide the model when faced with queries that share local features yet differ in specific causal semantics, thus compromising locality.
These limitations raise a critical question: How can knowledge editing orchestrate MLLM comprehension adaptively on cross-modal prompting, while balancing semantic transferability and robustness to spurious correlations?
@@
To this end, we rethink MLLM editing as an out-of-distribution (OOD) cross-modal semantic generalization problem. OOD (Ye et al., 2021; Montasser et al., 2024) refers to identifying invariant versus spurious features that drive distribution shifts, enabling the model to generalize to unseen domains. In Figure 1(b), editing MLLM involves partitioning causal scopes in prompts, which has two distributional shifts: (i) Semantic-shift indicates the shift from in-distribution editing scopes to neighboring regions, constituting generalized targets. Once the editing instills the mirror reflection principle into MLLM, the model can generalize from a narrowly edited instance (a tabby cat staring itself in the mirror) to similar scenarios (an orange cat or dog doing the same thing). (ii) Factual-shift denotes the transition from in-distribution to out-of-distribution regions, encompassing extraneous concepts beyond editing scopes. The mirror reflection principle should not be overapplied to counting prompts lacking mirror-specific visual features. Building on these, robust MLLM editing requires identifying invariant trajectories for cross-modal predictions and removing spurious factors that disrupt causal associations.
@@
In this paper, we propose a plug-and-play editing OOD optimization framework for multimodal LLM, termed ODEdit, which leverages cross-modal causal trajectory invariant learning to ensure knowledge editing robustness across diverse distributions. To explicitly enhance the MLLM’s discriminative awareness of semantic-shift and factual-shift, we first introduce a tripartite OOD risk formulation that imposes tailored constraints on in-distribution, semantic-neighboring, and out-of-distribution features. We apply the Kullback-Leibler divergence regularization to preserve locality while developing a maximum mean discrepancy-based metric learning to align representations of edited concepts and their semantic variants. To discern and stabilize the edit trajectories across heterogeneous cross-modal environments, we further propose Edit Trajectory Invariant Learning (ETIL). ETIL first reforms the editing OOD objective into an equivalent invariant risk minimization problem, where an environment-aware classifier is introduced to exploit feature invariance and irrelevance. Then, to suppress the sensitivity of the edit trajectory to spurious environmental changes, ETIL integrates a Total Variation factor as the penalty term in the risk estimation. The invariant risk minimization is achieved through a primal-dual optimization strategy, ensuring that the edited model captures reusable causal structures while filtering out superficial correlations.
@@ 2 Related Work
Multimodal LLM Editing. Recent advances in MLLMs (Li et al., 2023, 2024a; Ma et al., 2025; Luo et al., 2025b) have motivated research on multimodal knowledge editing (Pan et al., 2023; Zhou et al., ). A series of benchmarks, e.g., MMEdit (Cheng et al., 2023), MIKE (Li et al., 2024b), VLKEB (Huang et al., 2024), MC-MKE (Zhang et al., 2024a), MMKE (Du et al., 2025), provide unified datasets and evaluation to assess multimodal editing efficacy. However, research on strengthening the robustness of MLLM editing methods holistically across reliability, locality, and generality remains underexplored. MSCKE (Zeng et al., 2024) establishes a classifier-based knowledge editor to identify and update specific visual entities. UniKE (Pan et al., 2024) integrates intrinsic knowledge editing and external knowledge resorting to promote locality and generality. BalancEdit (Guo et al., 2025) performs codebook-based edits that balance generality and locality. Nevertheless, existing work remains constrained to rigid parameter-to-output mappings, which prevent MLLMs from intelligently distinguishing between semantic-shift and factual-shift, thereby hindering adaptive and robust editing.
Out-of-Distribution Generalization. OOD generalization is a core challenge in machine learning, aiming for generalization under covariate shift without access to data in the target domain (Muandet et al., 2013; Arjovsky et al., 2019; Zhou et al., 2025a). The mainstream works (Arjovsky et al., 2019; Krueger et al., 2021; Ahuja et al., 2020; Lai & Wang, 2024) utilize invariant risk minimization with regularizer to explore invariant representations across different training environments. Further, a wide range of techniques is leveraged to extract and generalize invariant features (Yu et al., 2023), e.g., context-based augmentation (Nam et al., 2021), representation alignment (Dou et al., 2019; Ruan et al., 2021), gradient manipulation (Shahtalebi et al., 2021), distributional robust optimization (Ghosal & Li, 2023), and meta-learning (Chen et al., 2023). In this paper, we make the first attempt to cast MLLM editing as an OOD generalization problem, where invariant learning across editing environments is enforced via a total invariance regularizer on cross-modal semantic features, so as to improve editing robustness and adaptability.
@@ 5.1 Experimental Setup
Baseline Methods. We describe four types of baselines and how we incorporate ODEdit into each method as a plug-and-play optimization framework in Appendix D.3.
@@ 5.4 In-Depth Analysis
Computational Cost. We pick one typical parameter-adjusting (MEND) and model-extending (WISE) baselines for comparison. From Appendix Table 6: 1) ODEdit introduces a supplementary network that processes parameters from the knowledge-editing layer, leading to increased memory usage. But this is the reasonable trade-off for the gains in performance and is acceptable given the current state of computational resources. 2) ODEdit does not incur a significant increase in time cost, indicating its efficiency. 3) While integrating ODEdit increases steps, the resultant increase in total time cost does not constitute an order-of-magnitude change and remains within a practical range for real-world deployment. Employing higher-performance computing resources would substantially reduce this training time gap.
@@ D.3 Baseline Methods
Models BLIP2-OPT MiniGPT-4
WISE 28.47GB 0.401 0.023 0.424 7 36.59GB 0.473 0.035 0.508 8
WISE+ODEdit 47.53GB 0.442 0.043 0.485 7 69.36GB 0.533 0.072 0.605 8
MEND 14.75GB 1.369 0.018 1.387 25000 25.30GB 1.676 0.188 1.865 10000

### v2

# Generalizable Multimodal Large Language Model Editing via Invariant Trajectory Learning

@@ Abstract
Knowledge editing emerges as a crucial technique for efficiently correcting incorrect or outdated knowledge in large language models (LLMs). Existing editing methods rely on a rigid mapping from parameter or module modifications to output, which causes the generalization limitation in Multimodal LLM (MLLM). In this paper, we reformulate MLLM editing as an out-of-distribution (OOD) generalization problem, where the goal is to discern semantic shift with factual shift and thus achieve robust editing among diverse cross-modal prompting. The key challenge of this OOD problem lies in identifying invariant causal trajectories that generalize accurately while suppressing spurious correlations. To address it, we propose ODEdit,a plug-and-play invariant learning based framework that optimizes the tripartite OOD risk objective to simultaneously enhance editing reliability, locality, and generality. We further introduce an edit trajectory invariant learning method, which integrates a total variation penalty into the risk minimization objective to stabilize edit trajectories against environmental variations. Theoretical analysis and extensive experiments demonstrate the effectiveness of ODEdit.
@@ 1 Introduction
With rapid applications of large language models (LLMs) (Liu et al., 2024), ensuring their knowledge correctness and currency in a cost-efficient manner has become a critical concern. Knowledge editing (De Cao et al., 2021; Wang et al., 2023, 2024b) is an emerging technique that supports data-efficient modifications on pre-trained models within a specific scope of knowledge. Existing LLM editing methods can be broadly categorized into: i) parameter-adjusting (Meng et al., 2022b; Fang et al., 2024; Jiang et al., 2025), which directly tune a subset of parameters in the original model, and ii) model-extending methods (Huang et al., 2023; Hartvigsen et al., 2023; Yu et al., 2024), which attach auxiliary components while keeping the backbone parameters intact. However, current studies remain largely confined to unimodal LLMs, leaving open their extension to multimodal LLMs (MLLMs) (Du et al., 2025; Guo et al., 2025).
Figure 1: The motivation of ODEdit. We illustrate how semantic-shift and factual-shift emerge in this editing OOD problem. We give out the formal definition of two shifts in Appendix C.
Several pioneer studies have made initial attempts to edit MLLMs. UniKE (Pan et al., 2024) performs both parameter-adjusting and model-extending, by representing intrinsic and external knowledge as separate vectorized key-value memories. BalancEdit (Guo et al., 2025) as a model-extending approach, generates positive and negative samples to determine the influence scope, and edits with a localized codebook in the latent space. Despite advances in promoting edit success (reliability), these works all operationalize editing as a rigid mapping from parameter modifications $\Delta W$ or auxiliary component $\Delta M$ to output variations $\Delta Y$, distilled from limited training samples. Nevertheless, within MLLM, the output variation is shaped by cascaded reasoning as unimodal perception → cross-modal alignment → shared semantic space reasoning, such that localized modification to parameters or modules can trigger complex and non-local semantic propagation effects. Relying on such a rigid mapping thus imposes the generalization problem from two aspects: i) First, it fails to disentangle the coherent shared semantic structures that span diverse cross-modal contexts. This precludes MLLM from discovering reusable semantic substrates, reducing editing to a brittle case-wise alignment exercise rather than principled semantic generality. ii) Second, it drives MLLM to memorize entrenched linkages between local features and outputs. The inflexible mapping misguides the model when faced with queries that share local features yet differ in specific causal semantics, thus compromising locality.
@@
To address it, we rethink MLLM editing as an out-of-distribution (OOD) cross-modal semantic generalization problem. OOD (Ye et al., 2021; Montasser et al., 2024) refers to identifying invariant versus spurious features that drive distribution shifts, enabling the model to generalize to unseen domains. In Figure 1, editing MLLM involves partitioning cross-modal semantic scopes in prompts, which has two distributional shifts: (i) Semantic-shift indicates the shift from in-distribution editing scopes to neighboring regions, constituting generalized targets. Once the editing instills the mirror reflection principle into MLLM, the model can generalize from a narrowly edited instance (a tabby cat staring itself in the mirror) to similar scenarios (an orange cat or dog doing the same thing). (ii) Factual-shift denotes the transition from in-distribution to out-of-distribution regions, encompassing extraneous concepts beyond editing scopes. The mirror reflection principle should not be overapplied to counting prompts when lacking mirror-specific visual features. Building on these, robust MLLM editing requires identifying invariant trajectories for cross-modal predictions and removing spurious factors that disrupt semantical reasoning associations.
@@
In this paper, we propose a plug-and-play editing OOD optimization framework for multimodal LLM, termed ODEdit, which leverages cross-modal reasoning trajectory invariant learning to ensure knowledge editing robustness across diverse distributions. To explicitly enhance the MLLM’s discriminative awareness of semantic-shift and factual-shift, we first introduce a tripartite OOD risk formulation that imposes tailored constraints on in-distribution, semantic-neighboring, and out-of-distribution features. We apply the Kullback-Leibler divergence regularization to preserve locality while developing a maximum mean discrepancy-based metric learning to align representations of edited concepts and their semantic variants. To discern and stabilize the edit trajectories across heterogeneous cross-modal environments, we further propose Edit Trajectory Invariant Learning (ETIL). ETIL first reforms the editing OOD objective into an equivalent invariant risk minimization problem, where an environment-aware classifier is introduced to exploit feature invariance and irrelevance. Then, to suppress the sensitivity of the edit trajectory to spurious environmental changes, ETIL integrates a Total Variation factor as the penalty term in the risk estimation. The invariant risk minimization is achieved through a primal-dual optimization strategy, ensuring that the edited model captures reusable causal structures while filtering out superficial correlations.
@@ 2 Related Work
Multimodal LLM Editing. Recent advances in MLLMs (Li et al., 2023, 2024a; Ma et al., 2025; Luo et al., 2025) have motivated research on multimodal knowledge editing (Pan et al., 2023; Zhou et al., ). A series of benchmarks, e.g., MMEdit (Cheng et al., 2023), MIKE (Li et al., 2024b), VLKEB (Huang et al., 2024), MC-MKE (Zhang et al., 2024a), MMKE (Du et al., 2025), provide unified datasets and evaluation to assess multimodal editing efficacy. However, research on strengthening the robustness of MLLM editing methods holistically across reliability, locality, and generality remains underexplored. MSCKE (Zeng et al., 2024) establishes a classifier-based knowledge editor to identify and update specific visual entities. UniKE (Pan et al., 2024) integrates intrinsic knowledge editing and external knowledge resorting to promote locality and generality. BalancEdit (Guo et al., 2025) performs codebook-based edits that balance generality and locality. Nevertheless, existing work remains constrained to rigid parameter-to-output mappings, which prevent MLLMs from intelligently distinguishing between semantic-shift and factual-shift, thereby hindering adaptive and robust editing.
Out-of-Distribution Generalization. OOD generalization is a core challenge in machine learning, aiming for generalization under covariate shift without access to data in the target domain (Muandet et al., 2013; Arjovsky et al., 2019; Zhou et al., 2025). The mainstream works (Arjovsky et al., 2019; Krueger et al., 2021; Ahuja et al., 2020; Lai & Wang, 2024) utilize invariant risk minimization with regularizer to explore invariant representations across different training environments. Further, a wide range of techniques is leveraged to extract and generalize invariant features (Yu et al., 2023), e.g., context-based augmentation (Nam et al., 2021), representation alignment (Dou et al., 2019; Ruan et al., 2021), gradient manipulation (Shahtalebi et al., 2021), distributional robust optimization (Ghosal & Li, 2023), and meta-learning (Chen et al., 2023). In this paper, we make the first attempt to cast MLLM editing as an OOD generalization problem, where invariant learning across editing environments is enforced via a total invariance regularizer on cross-modal semantic features, so as to improve editing robustness and adaptability.
@@ 5.1 Experimental Setup
Baseline Methods. We compare ODEdit with four types of baselines: (1) Fine-tuning: FT directly tunes the last three layers of MLLM. (2) Parameter-adjusting unimodal editing: MEND (Mitchell et al., 2021). (3) Model-extending unimodal editing: IKE (Zheng et al., 2023), SERAC (Mitchell et al., 2022), T-Patcher (Huang et al., 2023), WISE (Wang et al., 2024a). (4) Integrate parameter-adjusting and model-extending multimodal editing: UniKE (Pan et al., 2024). ODEdit serves as a plug-and-play universal framework, capable of being seamlessly integrated into any editing model that relies on loss-based optimization.
@@ 5.4 In-Depth Analysis
Models BLIP2-OPT MiniGPT-4
WISE 28.47GB 0.401 0.023 0.424 7 36.59GB 0.473 0.035 0.508 8
WISE+ODEdit 47.53GB 0.442 0.043 0.485 7 69.36GB 0.533 0.072 0.605 8
MEND 14.75GB 1.369 0.018 1.387 25000 25.30GB 1.676 0.188 1.865 10000
@@
Computational Cost. We pick one typical parameter-adjusting (MEND) and model-extending (WISE) baselines for comparison. From 6: 1) ODEdit introduces a supplementary network that processes parameters from the knowledge-editing layer, leading to increased memory usage. But this is the reasonable trade-off for the gains in performance and is acceptable given the current state of computational resources. 2) ODEdit does not incur a significant increase in time cost, indicating its efficiency. 3) While integrating ODEdit increases steps, the resultant increase in total time cost does not constitute an order-of-magnitude change and remains within a practical range for real-world deployment. Employing higher-performance computing resources can reduce this training time gap.